# Deblur4DGS: 4D Gaussian Splatting from Blurry Monocular Video

### v1

@@ Abstract
Recent 4D reconstruction methods have yielded impressive results but rely on sharp videos as supervision. However, motion blur often occurs in videos due to camera shake and object movement, while existing methods render blurry results when using such videos for reconstructing 4D models. Although a few NeRF-based approaches attempted to address the problem, they struggled to produce high-quality results, due to the inaccuracy in estimating continuous dynamic representations within the exposure time. Encouraged by recent works in 3D motion trajectory modeling using 3D Gaussian Splatting (3DGS), we suggest taking 3DGS as the scene representation manner, and propose the first 4D Gaussian Splatting framework to reconstruct a high-quality 4D model from blurry monocular video, named Deblur4DGS. Specifically, we transform continuous dynamic representations estimation within an exposure time into the exposure time estimation. Moreover, we introduce exposure regularization to avoid trivial solutions, as well as multi-frame and multi-resolution consistency ones to alleviate artifacts. Furthermore, to better represent objects with large motion, we suggest blur-aware variable canonical Gaussians. Beyond novel-view synthesis, Deblur4DGS can be applied to improve blurry video from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive experiments on the above four tasks show that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods. The codes are available at https://github.com/ZcsrenlongZ/Deblur4DGS.
1 Introduction
Figure 1: Performance and rendering speed comparisons for novel-view synthesis on the 720 $\times$ 1080 images. Deblur4DGS produces better results with real-time rendering speed.
@@ 1 Introduction
Substantial efforts have been made for 4D reconstruction, which has extensive applications in augmented reality and virtual reality. To model static scenes, Neural Radiance Field (NeRF) [52] and 3D Gaussian Splatting (3DGS) [27] propose implicit neural representation manner and explicit Gaussian ellipsoids one, respectively. To model dynamic objects, implicit neural fields [48, 108, 90, 84, 5, 60, 87] and explicit deformation [10, 8, 49, 89, 25, 30, 44, 40, 79] are suggested for motion representation. While achieving great progress, most methods rely on synchronized multi-view videos. They yield unsatisfactory results when applied to monocular video, where dynamic objects are only observed once at each timestamp. To alleviate the under-constrained nature of the problem, recent studies have introduced data-driven priors, such as depth maps [35, 88], optical flows [16, 76, 17, 19, 46, 108, 35, 80], tracks [79, 65, 70, 36], and generative models [86, 8, 91, 93] to improve 4D reconstruction.
@@
Unfortunately, motion blur often arises in the videos due to camera shake and object movement, while the above methods heavily rely on sharp videos for supervision. When reconstructing the 4D scene from the blurry video, these methods can render blurry results. The first step to solving this problem is to deal with camera motion blur, which is relatively simple. Some NeRF-based [32, 50, 78, 33, 39] and 3DGS-based [102, 58, 7, 55] methods have suggested jointly optimizing 3D static scene representation and camera poses within the exposure time by calculating the reconstruction loss between the synthetic blurry images and the input blurry frames. In contrast, the object motion blur is more challenging to address, as the solution has to estimate continuous and sharp dynamic representations within the exposure time to simulate blurry frames. As far as we know, only DybluRF [72, 4] attempted this with NeRF but struggled to produce high-quality and real-time rendering results, as shown in Fig. 1.
@@
In this work, we propose taking 3DGS [27] as the scene representation manner to explore the problem, driven by two main motivations. First, the real-time rendering capability of 3DGS and its successful application in 4D reconstruction make this method highly promising. Second, the explicit representation property of 3DGS makes it easier to model 3D object motion than NeRF. In particular, recent advances [30, 79, 65, 70] in reconstructing 3D motion trajectories present an opportunity to simplify this continuous dynamic representations estimation problem. Based on this, the challenging problem can be transformed into simple exposure time estimation, without the extra motion trajectory modeling in DybluRF [72, 4]. With the estimated exposure time, continuous dynamic representations can be obtained by directly interpolating between representations at the nearest integer timestamps.
@@
Specifically, we name it Deblur4DGS, which is the first Gaussian Splatting framework designed for 4D reconstruction from blurry monocular video. For the static scene, we present a camera motion predictor to learn camera poses at exposure start and end, which is jointly optimized with static Gaussians. For the dynamic objects, we optimize learnable exposure time parameters and dynamic Gaussians of the integer timestamps, simultaneously. Then continuous camera poses and dynamic Gaussians within exposure time can be obtained by interpolation, and they are used to render continuous sharp frames to calculate the reconstruction loss. Moreover, the exposure regularization term is introduced to avoid trivial solutions, while multi-frame and multi-resolution consistency regularization terms are used to alleviate artifacts. Furthermore, existing 4D reconstruction methods generally select Gaussians at a single timestamp as canonical Gaussians (e.g., the ones at the first timestamp). However, it may produce results with missing details in scenes with large motion, especially when processing blurry videos with a low frame rate. To alleviate this issue, we suggest variable canonical Gaussians as time progresses based on the image blur level. Gaussians corresponding to the sharper frame are selected as the canonical ones for better blur removal, and each canonical Gaussian is only used for some nearby timestamps to reduce the difficulty of modeling large motion.
@@
Blurry videos suffer from not only motion blur, but also low frame rates and scene shake generally. Beyond novel-view synthesis, the optimized Deblur4DGS can also be applied to address these problems, achieving deblurring, frame interpolation, and video stabilization. We evaluate the 4D reconstruction quality of Deblur4DGS from all four perspectives. Extensive experiments demonstrate that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods quantitatively and qualitatively while maintaining real-time rendering speed. Furthermore, Deblur4DGS has competitive capabilities in comparison with task-specific video processing models trained in a supervised manner.
@@
  • We propose Deblur4DGS, the first 4D Gaussian Splatting framework specially designed to reconstruct a high-quality 4D model from blurry monocular video.
@@
  • We propose transforming dynamic representation estimation into exposure time estimation, where a series of regularizations are suggested to tackle the under-constrained optimization and blur-aware variable canonical Gaussians is present to better represent objects with large motion.
@@
  • Extensive experiments on novel-view synthesis, deblurring, frame interpolation, and video stabilization tasks show that Deblur4DGS significantly outperforms state-of-the-art 4D reconstruction methods.
2 Related Work
2.1 Image and Video Deblurring
@@
Deep learning-based image [73, 62, 38, 29, 82, 92, 37, 99, 43, 101] and video [23, 54, 69, 83, 106, 42, 41, 56, 95, 66, 105, 24, 104, 6, 96] deblurring methods have been widely explored. Compared to image deblurring methods, video ones can leverage temporal clues between consecutive frames for more effective restoration. For example, DBLRNet [96] employs spatial-temporal 3D convolution for spatial and temporal features aggregation. DSTNet [56] develops a deep discriminative spatial and temporal network. BasicVSR++ [6] improves feature fusion with second-order feature propagation and flow-guided deformable alignment. BSSTNet [95] introduces a blur map to sufficiently utilize the entire video, achieving recent state-of-the-art. When reconstructing from a blurry video, pre-processing it with the 2D deblurring method is a straightforward manner. However, 2D deblurring methods cannot perceive 3D structures and maintain scene geometric consistency, leading to unsatisfactory scene reconstruction.
@@
Figure 2: (a) Training of Deblur4DGS. When processing $t$-th frame, we first discretize its exposure time into $N$ timestamps. Then, we estimate continuous camera poses $\{\mathbf{P}_{t,i}\}_{i=1}^{N}$ (see Sec. 4.1) and dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ (see Sec. 4.2) within exposure time. Next, we render each latent sharp image $\hat{\mathbf{I}}_{t,i}$ with the camera pose $\mathbf{P}_{t,i}$, dynamic Gaussians $\mathbf{D}_{t,i}$ and static Gaussians $\mathbf{S}$. Finally, $\{\hat{\mathbf{I}}_{t,i}\}_{i=1}^{N}$ are averaged to obtain the synthetic blurry image $\hat{\mathbf{B}}_{t}$, which is used to calculate the reconstruction loss $\mathcal{L}_{rec}$ with the given blurry frame $\mathbf{B}_{t}$. To regularize the under-constrained optimization, we introduce exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$ and multi-resolution consistency regularization $\mathcal{L}_{mrc}$ (see Sec. 4.3). (b) Rendering of Deblur4DGS. Deblur4DGS produces the sharp image with user-provided timestamp $t$ and camera pose $\mathbf{P}_{t}$.
2.2 3D and 4D Reconstruction
@@
To reconstruct 3D models, NeRF [52] and 3DGS [27] introduce implicit neural representation manner and explicit Gaussian ellipsoids one respectively, where the latter generally achieves better results in both image quality and rendering speed. To reconstruct 4D models, most works [60, 87, 5, 15, 13, 45, 18, 103, 14, 68, 11, 48, 49, 89, 30, 20, 44, 40, 67, 74, 84, 90, 1, 51, 10, 8, 25] incorporate implicit neural fields and explicit deformation for motion representation. For example, DeformableGS [90] achieves it with multi-layer perceptrons (MLPs). 4DGasusians [1] improves it with a multi-resolution HexPlane [5]. E-D3DGS [84] introduces per-Gaussian embeddings and temporal embeddings. Moreover, to better reconstruct from monocular video, some studies enhance 4D reconstruction with data-driven priors, such as depth maps [35, 88], optical flows [16, 76, 17, 19, 46, 108, 35, 80], tracks [79, 65, 70, 36], and generative models [86, 8, 91, 93, 98].
@@
Note that these methods heavily rely on high-quality sharp videos for supervision and perform poorly when facing blurry inputs. To process camera motion in static areas, recent works [32, 50, 78, 33, 39, 31, 102, 58, 34, 7, 55] suggest jointly optimizing the scene representation and recovering the camera poses within the exposure time. To process object motion blur in dynamic scenes, DybluRF [72, 4] incorporates object motion blur formation into dynamic model optimization but faces challenges in producing high-quality images and achieving real-time rendering. In this work, with 3DGS [27] as the scene representation manner, we develop Deblur4DGS to reconstruct a high-quality 4D model from a blurry video.
3 Preliminary
3.1 4D Gaussian Splatting
@@ 3 Preliminary
A 3D Gaussian [27] is parameterized by $\{\mathbf{x},\mathbf{r},\mathbf{s},\mathbf{o},\mathbf{c}\}$, where $\mathbf{x}$ characterizes the center position in the world space, rotation matrix $\mathbf{r}$ and scale matrix $\mathbf{s}$ define the shape, $\alpha$ is opacity, and spherical harmonics (SH) coefficients $\mathbf{c}$ represent the view-dependent color.
@@
In 4D Gaussian Splatting (4DGS), the static and dynamic regions are usually processed, respectively. Static regions can be represented by a set of 3D Gaussians, named $\mathbf{S}$. For the dynamic areas, 4DGS generally selects a timestamp (e.g., the first timestamp) and represents the objects by canonical dynamic Gaussians, i.e., $\mathbf{C}$. Then, $\mathbf{C}$ is deformed to other timestamps by an implicit neural field [48, 108, 90, 84, 5, 60, 87] or explicit manner [10, 8, 49, 89, 25, 30, 44, 40, 79] for motion representation. Denote by $\mathbf{D}_{t}$ the dynamic Gaussians at $t$-th timestamp, it can be written as,
$\mathbf{D}_{t}=\mathcal{F}(\mathbf{C},t;\Theta_{\mathcal{F}}).\vspace{-2mm}$ (1)
@@
3.2 Motion Blur Formation
Motion blur occurs due to camera shake and object movement, which can be simply regarded as the integration of a latent sharp image sequence [53], i.e.,
[…]
$\mathbf{B}(u,v)\in\mathbb{R}^{H\times W\times 3}$ is the blurry image and $\mathbf{I}_{t}(u,v)$ is the latent sharp one at $t$-th timestamp. $(u,v)$ is pixel location, $\tau$ is the camera exposure time, and $\phi$ is a normalization factor. To approximate the integral operation, recent works [102, 72, 78] divide the exposure time into $N$ timestamps and regard the blurry image as the average of $N$ sharp images, i.e.,
$\mathbf{B}(u,v)\approx\frac{1}{N}\sum_{i=0}^{N-1}\mathbf{I}_{i}(u,v).\vspace{-2mm}$ (3)
In this work, we reconstruct a high-quality 4D model from a blurry video by incorporating the physical motion blur formation into model optimization.
4 Proposed Method
Let $\{\mathbf{B}_{t}\}_{t=1}^{T}$ and $\{\mathbf{M}_{t}\}_{t=1}^{T}$ denote a blurry video with $T$ timestamps and the corresponding masks indicating dynamic areas, respectively. As shown in Fig. 2(a), when processing $t$-th frame, we first evenly divide its camera exposure time into $N$ timestamps. Then, we estimate continuous camera poses $\{\mathbf{P}_{t,i}\}_{i=1}^{N}$ and dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ to simulate camera shake and object movement. Next, we render each sharp image $\hat{\mathbf{I}}_{t,i}$ with the corresponding camera pose $\mathbf{P}_{t,i}$, dynamic Gaussians $\mathbf{D}_{t,i}$ and static Gaussians $\mathbf{S}$. After that, we average $\{\hat{\mathbf{I}}_{t,i}\}_{i=1}^{N}$ to obtain the synthetic blurry image $\hat{\mathbf{B}}_{t}$, which is used to calculate the reconstruction loss $\mathcal{L}_{rec}$ with the given blurry frame $\mathbf{B}_{t}$, i.e.,
$\mathcal{L}_{rec}=(1-\beta)\mathcal{L}_{1}(\hat{\mathbf{B}}_{t},\mathbf{B}_{t})+\beta\mathcal{L}_{ssim}(\hat{\mathbf{B}}_{t},\mathbf{B}_{t}).\vspace{-1mm}$ (4)
$\mathcal{L}_{1}$ and $\mathcal{L}_{ssim}$ are $\ell_{1}$ loss and SSIM [81] loss, respectively. $\beta$ is set to $0.2$. The settings all follow 3DGS [27].
4.1 Continuous Camera Poses Estimation
To estimate continuous camera poses, recent methods [102, 58, 7, 55] directly optimize exposure start and end poses (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$). Then, the linear interpolation is performed between $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$ in the Lie-algebra of $\mathbf{SE}(3)$ to obtain the camera pose at $i$-th intermediate timestamp (i.e., $\mathbf{P}_{t,i}$). It can be written as,
$\mathbf{P}_{t,i}=\mathbf{P}_{t,1}\odot\texttt{exp}(\frac{i-1}{N-1}\odot\texttt{log}(\frac{\mathbf{P}_{t,N}}{\mathbf{P}_{t,1}})).\vspace{-2mm}$ (5)
@@ 4.1 Continuous Camera Poses Estimation
We follow the manner but deploy a tiny MLP as the camera motion predictor (see details in Suppl.) for more stable optimization. We pre-train the camera motion predictor and static Gaussians $\mathbf{S}$ with static reconstruction loss $\mathcal{L}_{rec}^{s}$, i.e.,
$\mathcal{L}_{rec}^{s}=(1-\beta)\mathcal{L}_{1}(\hat{\mathbf{B}}_{t}^{s},\mathbf{B}_{t}^{s})+\beta\mathcal{L}_{ssim}(\hat{\mathbf{B}}_{t}^{s},\mathbf{B}_{t}^{s}).\vspace{-1mm}$ (6)
@@
4.2 Continuous Dynamic Gaussians Estimation

### v2

@@ Abstract
Recent 4D reconstruction methods have yielded impressive results but rely on sharp videos as supervision. However, motion blur often occurs in videos due to camera shake and object movement, while existing methods render blurry results when using such videos for reconstructing 4D models. Although a few approaches attempted to address the problem, they struggled to produce high-quality results, due to the inaccuracy in estimating continuous dynamic representations within the exposure time. Encouraged by recent works in 3D motion trajectory modeling using 3D Gaussian Splatting (3DGS), we take 3DGS as the scene representation manner, and propose Deblur4DGS to reconstruct a high-quality 4D model from blurry monocular video. Specifically, we transform continuous dynamic representations estimation within an exposure time into the exposure time estimation. Moreover, we introduce the exposure regularization term, multi-frame, and multi-resolution consistency regularization term to avoid trivial solutions. Furthermore, to better represent objects with large motion, we suggest blur-aware variable canonical Gaussians. Beyond novel-view synthesis, Deblur4DGS can be applied to improve blurry video from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive experiments in both synthetic and real-world data on the above four tasks show that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods. The codes will be publicly available.
Introduction
@@
Substantial efforts have been made for 4D reconstruction, which has extensive applications in augmented reality and virtual reality. To model static scenes, Neural Radiance Field (NeRF) (Mildenhall et al. 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) propose implicit neural representation manner and explicit Gaussian ellipsoids one, respectively. To model dynamic objects, implicit neural fields (Zhu et al. 2024; Yang et al. 2024; Wu et al. 2024; Yan, Li, and Lee 2023) and explicit deformation (Duan et al. 2024b; Chu, Ke, and Fragkiadaki 2024; Katsumata, Vo, and Nakayama 2024; Lin et al. 2024b; Li et al. 2024; Wang et al. 2024) are suggested for motion representation. While achieving great progress, most methods rely on synchronized multi-view videos. They yield unsatisfactory results when applied to monocular video, where dynamic objects are only observed once at each timestamp. To alleviate the under-constrained nature of the problem, recent studies have introduced data-driven priors, such as depth maps (Lee et al. 2023c; Yang et al. ), optical flows (Gao et al. 2024; Zhu et al. 2024), tracks (Seidenschwarz et al. 2024; Lei et al. 2024), and generative models (Wu et al. 2025; Chu, Ke, and Fragkiadaki 2024; Zeng et al. 2025) for better 4D reconstruction.
@@
Unfortunately, motion blur often arises due to camera shake and object movement. When reconstructing the 4D scene from the blurry video, the above methods usually render blurry results. The first step to solving this problem is to deal with camera motion blur, which is relatively simple. Some NeRF-based (Lee et al. 2023a; Wang et al. 2023; Lee et al. 2023b) and 3DGS-based (Zhao, Wang, and Liu 2024; Chen and Liu 2024; Oh et al. 2024) methods have suggested jointly optimizing 3D representation and camera poses within the exposure time by calculating the reconstruction loss between the synthetic blurry images and the input blurry frames. In contrast, the object motion blur is more challenging to address, as the solution has to estimate continuous and sharp dynamic representations within the exposure time to simulate blurry frames.
@@
In this work, we take 3DGS (Kerbl et al. 2023) as the scene representation manner to explore the problem, driven by two main motivations. First, its successful application in 4D reconstruction make this method highly promising. Second, the explicit 3D motion modeling presents an opportunity to simplify the complex continuous dynamic representations estimation within the exposure time into exposure time estimation, avoiding complex extra motion modeling in DyBluRF (Sun et al. 2024a; Bui et al. 2023). Once the exposure time is estimated, continuous dynamic representations can be obtained by directly interpolating between representations at the nearest integer timestamps. We note that the concurrent work BARD-GS (Lu et al. 2025) adopts a similar strategy, but they perform unsatisfactorily due to the under-constrained optimization, especially for large object motion.
@@
Specifically, we propose Deblur4DGS, a Gaussian Splatting framework for 4D reconstruction from blurry monocular video. For the static scene, we jointly optimize the camera poses at exposure start and end with static Gaussians. For the dynamic objects, we optimize learnable exposure time parameters and dynamic Gaussians of the integer timestamps, simultaneously. Then continuous camera poses and dynamic Gaussians within exposure time can be obtained by interpolation, and they are used to render continuous sharp frames to calculate the reconstruction loss. Moreover, to avoid trivial solutions, we introduce the exposure regularization term, as well as the multi-frame and multi-resolution consistency regularization terms. Furthermore, existing 4D reconstruction methods generally select Gaussians at a single timestamp as canonical Gaussians. However, it may produce results with missing details in scenes with large motion, especially when processing blurry videos with a low frame rate. To alleviate this issue, we suggest variable canonical Gaussians as time progresses based on the image blur level. Gaussians corresponding to the sharper frame are selected as the canonical ones for better blur removal, and each canonical Gaussian is only used for nearby timestamps to reduce difficulty of modeling large motion.
@@
Blurry videos suffer from not only motion blur, but also low frame rates and scene shake generally. Beyond novel-view synthesis, the optimized Deblur4DGS can be applied to address these problems, achieving deblurring, frame interpolation, and video stabilization. We evaluate Deblur4DGS from all four perspectives. Extensive experiments on both synthetic and real-world data demonstrate that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods quantitatively and qualitatively while maintaining real-time rendering speed. Furthermore, Deblur4DGS has competitive capabilities in comparison with task-specific video processing models trained in a supervised manner.
@@
  • We propose Deblur4DGS, a 4D Gaussian Splatting framework specially designed to reconstruct a high-quality 4D model from blurry monocular video.
@@
  • We propose transforming dynamic representation estimation into exposure time estimation, where a series of regularizations are suggested to tackle under-constrained optimization and blur-aware variable canonical Gaussians is present to better represent dynamic objects.
@@
  • Extensive experiments in synthetic and real-world data show that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods on novel-view synthesis, deblurring, frame interpolation, and video stabilization tasks.
Related Work
Image and Video Deblurring
@@
Deep learning-based image (Ren et al. 2023; Li et al. 2023; Kong et al. 2023; Wang et al. 2022; Zhang et al. 2022; Lin et al. 2024a; Zhang et al. 2024) and video (Zhong et al. 2023b; Liang et al. 2024; Pan et al. 2023; Zhang, Xie, and Yao 2024; Zhong et al. 2023a, 2020; Chan et al. 2022) deblurring methods have been widely explored. Compared to image deblurring methods, video ones leverage temporal clues between consecutive frames for more effective restoration. DBLRNet (Zhang et al. 2018a) employs spatial-temporal 3D convolution for spatial and temporal features aggregation. DSTNet (Pan et al. 2023) develops a deep discriminative spatial and temporal network. BasicVSR++ (Chan et al. 2022) improves feature fusion with second-order feature propagation and flow-guided alignment. BSSTNet (Zhang, Xie, and Yao 2024) introduces a blur map to sufficiently utilize the entire video, achieving recent state-of-the-art. When reconstructing from a blurry video, pre-processing it with the 2D deblurring method is a straightforward manner. However, 2D deblurring methods cannot perceive 3D structures and maintain scene geometric consistency, leading to unsatisfactory scene reconstruction.
@@
Figure 1: (a) Training of Deblur4DGS. When processing $t$-th frame, we first discretize its exposure time into $N$ timestamps. Then, we estimate continuous camera poses $\{\mathbf{P}_{t,i}\}_{i=1}^{N}$ and dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ within exposure time. Next, we render each latent sharp image $\hat{\mathbf{I}}_{t,i}$ with the camera pose $\mathbf{P}_{t,i}$, dynamic Gaussians $\mathbf{D}_{t,i}$ and static Gaussians $\mathbf{S}$. Finally, $\{\hat{\mathbf{I}}_{t,i}\}_{i=1}^{N}$ are averaged to obtain the synthetic blurry image $\hat{\mathbf{B}}_{t}$, which is used to calculate the reconstruction loss $\mathcal{L}_{rec}$ with the given blurry frame $\mathbf{B}_{t}$. To regularize the under-constrained optimization, we introduce exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$ and multi-resolution consistency regularization $\mathcal{L}_{mrc}$. (b) Rendering of Deblur4DGS. Deblur4DGS produces the sharp image with user-provided timestamp $t$ and camera pose $\mathbf{P}_{t}$.
3D and 4D Reconstruction
@@
To reconstruct 3D models, NeRF (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) introduce implicit neural representation manner and explicit Gaussian ellipsoids one respectively, where the latter generally achieves better results in both image quality and rendering speed. To reconstruct 4D models, most works (Somraj et al. 2024; Duan et al. 2024a; Lu et al. 2024; Luiten et al. 2023; Kratimenos, Lei, and Daniilidis 2023; Huang et al. 2024; Lin et al. 2024b; Li et al. 2024; Shaw et al. ; Sun et al. 2024b; Wu et al. 2024; Yang et al. 2024; Bae et al. 2024; Mihajlovic et al. 2024; Duan et al. 2024b) incorporate implicit neural fields and explicit deformation for motion representation. For example, DeformableGS (Yang et al. 2024) achieves it with multi-layer perceptrons. 4DGasusians (Bae et al. 2024) improves it with a multi-resolution HexPlane (Cao and Johnson 2023). E-D3DGS (Wu et al. 2024) introduces per-Gaussian embeddings and temporal embeddings. Moreover, to better reconstruct from monocular video, some studies enhance 4D reconstruction with data-driven priors, such as depth maps (Lee et al. 2023c; Yang et al. ), optical flows (Tian, Du, and Duan 2023; Gao et al. 2024), tracks (Wang et al. 2024; Seidenschwarz et al. 2024), and generative models (Wu et al. 2025; Chu, Ke, and Fragkiadaki 2024).
@@
Note that these methods heavily rely on high-quality sharp videos for supervision and perform poorly when facing blurry inputs. To process camera motion in static areas, recent works (Lee et al. 2023a; Ma et al. 2022; Wang et al. 2023; Lee et al. 2023b, 2024a; Zhao, Wang, and Liu 2024; Peng et al. 2024a; Lee et al. 2024b; Chen and Liu 2024; Oh et al. 2024) suggest jointly optimizing the scene representation and recovering the camera poses within the exposure time. To process object motion blur in dynamic scenes, DyBluRF (Sun et al. 2024a; Bui et al. 2023) incorporates object motion blur formation into dynamic model optimization but faces challenges in producing high-quality images and achieving real-time rendering. In this work, with 3DGS as the scene representation manner, we develop Deblur4DGS to reconstruct a high-quality 4D model from a blurry video.
Preliminary
4D Gaussian Splatting
@@
A 3D Gaussian (Kerbl et al. 2023) is parameterized by $\{\mathbf{x},\mathbf{r},\mathbf{s},\mathbf{o},\mathbf{c}\}$, where $\mathbf{x}$ characterizes the center position in the world space, rotation matrix $\mathbf{r}$ and scale matrix $\mathbf{s}$ define the shape, $\alpha$ is opacity, and spherical harmonics (SH) coefficients $\mathbf{c}$ represent the view-dependent color.
@@
4D Gaussian Splatting (4DGS) usually process static and dynamic regions separately. Static regions can be represented by a set of 3D Gaussians, named $\mathbf{S}$. For the dynamic areas, 4DGS generally selects a timestamp (e.g., the first timestamp) and represents the objects by canonical dynamic Gaussians, i.e., $\mathbf{C}$. Then, $\mathbf{C}$ is deformed to other timestamps for motion representation. Denote by $\mathbf{D}_{t}$ the dynamic Gaussians at $t$-th timestamp, it can be written as,
$\mathbf{D}_{t}=\mathcal{F}(\mathbf{C},t;\Theta_{\mathcal{F}}).$ (1)
@@
Motion Blur Formation
Motion blur occurs due to camera shake and object movement, which can be regarded as the integration of latent sharp images (Nah, Hyun Kim, and Mu Lee 2017), i.e.,
$\mathbf{B}(u,v)=\phi\int_{0}^{\tau}\mathbf{I}_{t}(u,v)dt.$ (2)
$\mathbf{B}\in\mathbb{R}^{H\times W\times 3}$ is the blurry image and $\mathbf{I}_{t}$ is the latent sharp one at $t$-th timestamp. $(u,v)$ is pixel location, $\tau$ is the camera exposure time, and $\phi$ is a normalization factor. To approximate the integral operation, recent works (Zhao, Wang, and Liu 2024; Sun et al. 2024a; Wang et al. 2023) divide the exposure time into $N$ timestamps and regard the blurry image as the average of $N$ sharp images, i.e.,
$\mathbf{B}(u,v)\approx\frac{1}{N}\sum_{i=0}^{N-1}\mathbf{I}_{i}(u,v).$ (3)
In this work, we reconstruct 4D model from a blurry video by integrating the blur formation into model optimization.
Proposed Method
Let $\{\mathbf{B}_{t}\}_{t=1}^{T}$ and $\{\mathbf{M}_{t}\}_{t=1}^{T}$ denote a blurry video with $T$ timestamps and the corresponding masks indicating dynamic areas (extracted by SAM2 (Ravi et al. 2024)), respectively. As shown in fig. 1(a), when processing $t$-th frame, we first evenly divide its camera exposure time into $N$ timestamps. Then, we estimate continuous camera poses $\{\mathbf{P}_{t,i}\}_{i=1}^{N}$ and dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ to simulate camera shake and object movement. Next, we render each sharp image $\hat{\mathbf{I}}_{t,i}$ with the corresponding camera pose $\mathbf{P}_{t,i}$, dynamic Gaussians $\mathbf{D}_{t,i}$ and static Gaussians $\mathbf{S}$. After that, we average $\{\hat{\mathbf{I}}_{t,i}\}_{i=1}^{N}$ to obtain the synthetic blurry image $\hat{\mathbf{B}}_{t}$, which is used to calculate the reconstruction loss $\mathcal{L}_{rec}$ with the given blurry frame $\mathbf{B}_{t}$, i.e.,
$\mathcal{L}_{rec}=(1-\beta)\mathcal{L}_{1}(\hat{\mathbf{B}}_{t},\mathbf{B}_{t})+\beta\mathcal{L}_{ssim}(\hat{\mathbf{B}}_{t},\mathbf{B}_{t}).$ (4)
$\mathcal{L}_{1}$ and $\mathcal{L}_{ssim}$ are $\ell_{1}$ loss and SSIM (Wang et al. 2004) loss, respectively. $\beta$ is set to $0.2$. The setting all follows 3DGS (Kerbl et al. 2023).
Continuous Camera Poses Estimation
To estimate continuous camera poses, recent methods (Zhao, Wang, and Liu 2024; Peng et al. 2024a; Chen and Liu 2024; Oh et al. 2024) directly optimize exposure start and end poses (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$). Then, the linear interpolation is performed between $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$ to obtain the camera pose at $i$-th intermediate timestamp (i.e., $\mathbf{P}_{t,i}$),i.e.,
$\mathbf{P}_{t,i}=\mathbf{P}_{t,1}\odot\texttt{exp}(\frac{i-1}{N-1}\odot\texttt{log}(\frac{\mathbf{P}_{t,N}}{\mathbf{P}_{t,1}})).$ (5)
@@
We follow the manner but deploy a tiny MLP as the camera motion predictor (see details in the Sec.A of Suppl) for more stable optimization. We pre-train it and static Gaussians $\mathbf{S}$ with static reconstruction loss $\mathcal{L}_{rec}^{s}$, i.e.,
$\mathcal{L}_{rec}^{s}=(1-\beta)\mathcal{L}_{1}(\hat{\mathbf{B}}_{t}^{s},\mathbf{B}_{t}^{s})+\beta\mathcal{L}_{ssim}(\hat{\mathbf{B}}_{t}^{s},\mathbf{B}_{t}^{s}).$ (6)
@@
Continuous Dynamic Gaussians Estimation